OpenStudy (anonymous):

SIMPLIFY 34/3 (FRACTION)

OpenStudy (anonymous):

11 1/3

OpenStudy (anonymous):

*as a mixed number

OpenStudy (anonymous):

so if i want to simplify 35/7 would it just be 5

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Because a fraction is the same as dividing.

hero (hero):

Simplification steps for 34/3 \[\frac{34}{3} = \frac{33 + 1}{3} = \frac{33}{3} + \frac{1}{3} = 11 + \frac{1}{3}\]

OpenStudy (anonymous):

all right thank you i think i got it now

hero (hero):

It is important to understand the difference between how to simplify a proper fraction and how to simplify an improper fraction.

OpenStudy (anonymous):

yeah I'm stuck on trying to to figure out how to reduce a improper fraction, and adding or subtracting fractions.

hero (hero):

Suppose you had to simplify \[\frac{34}{13}\] How do you think it should be done?

OpenStudy (anonymous):

well i would see how many times can 13 go into 34 and it will only go 2 times so i would put 2 8/13

hero (hero):

Yes, and to demonstrate this, you could proceed as follows: \[\frac{34}{13} = \frac{26 + 8}{13} = \frac{26}{13} + \frac{8}{13} = 2 \frac{8}{13}\]

OpenStudy (anonymous):

oh okay got it sorry not trying to bug but can you please help me on trying to understand how to add, subtract, divide, & multiply fractions for example. 5 7/8 + 2 3/4= 3 2/9 - 8/11= 2/7 x 15= 2 6/7 % 4 2/3= (if you can't its fine sorry this is too much)

hero (hero):

Well, to be honest, I don't think you will understand it all in one gulp. I believe the best thing for you in this case is to consult khan academy videos.

OpenStudy (anonymous):

oh okay thank you I'm just trying to study for a test on monday but as you can tell I'm not that good in fractions